OpenStudy (anonymous):

The price of silver has increased by 15% per year from 2000. In the year 2000, Harry bought a silver bowl for $190. Which of the following functions f(x) can be used to represent the price of the bowl x years after 2000?

OpenStudy (anonymous):

A. f(x) = 190(0.85)x B. f(x) = 190(1.15)x C. f(x) = 15(190)x D. f(x) = 15(0.85)x

OpenStudy (amistre64):

which one makes the most sense to you?

OpenStudy (anonymous):

After year 1 its 190x1.15 After year 2 its (190x1.15)x1.15 etc. So its there's a power missing in one of the a/b/c/d answers. I imagine it is \[190 \times 1.15^x\]

OpenStudy (anonymous):

thank you!!